1/2 times y divided by 3 equals 2
how do i figure out what y is

Answer by ptaylor(2198)

Well, we know that 1/2 times y is y/2 just like we know that 1/2 times 4 is 4/2=2
Now if we divide that by 3, we have (y/2)/3. This is a complex fraction and we can simplify it by multiplying the numerator and denominator by (1/3), like so:
(y/2)(1/3)/3(1/3) and this reduces to (y/6)/1 or y/6
And we are told that this equals 2, sooooo
(y/6)=2 multiply each side by 6
y=12
CK
1/2 times 12=6; 6 divided by 3=2
Hope this helps---ptaylor